SPECTROSCOPY OF THORIUM MONOXIDE, ThO; E(O^{+}),F(O^{+}),-X^{1}Σ^{+} BANDS

\underline{FANG WANG} AND TIMOTHY C. STEIMLE , Department of Chemistry and Biochemistry, Arizona State University, Tempe,AZ 85287; MICHAEL HEAVEN, Department of Chemistry, Emory University, Atlanta,GA 30322.

Thorium monoxide, ThO, has recently attracted interest as a possible venue for the determination of the electric dipole moment of the electron, d_{e}^{a}. Here we report on the results of an optical Stark study of the E(O^{+})-X^{1}Σ^{+}(1,0) band and the field-free study of the F (O^{+})-X^{1}Σ^{+}(0,0) band^{b}. A supersonic molecular beam of ThO was generated using a laser ablation technique and probed using laser excitation spectroscopy. The determined values for the permanent electric dipole moments, μ_{el}, for the E(O^{+})(v = 1) and X^{1}Σ^{+}(v = 0) vibronic states were determined to be 3.534±0.010 D and 2.782±0.012 D, respectively ^{c}. The dispersed laser induced fluorescence resulting from the excitation of the E(O^{+})- X^{1}Σ^{+}(1,0) and F (O^{+})-X^{1}Σ^{+}(0,0) bands have been recorded and the results are compared to Franck-Condon predictions. The radiative lifetimes for the E(O^{+})-X^{1}Σ^{+}(1,0) band F (O^{+})-X^{1}Σ^{+}(0,0) bands were determined.

aA. C. Vutha, W. C. Campbell, Y. V. Gurevich, N. R. Hutzler, M. Parsons, D. Patterson, E. Petrik, B. Spaun, J. M. Doyle, G.

Gabrielse, and D. DeMille, J. Phys. B: At., Mol. Opt. Phys.,43 074007/1 2010.

bG. Edvinsson and A. Lagerqvist, Physica Scripta, 32 602 1985.

cF. Wang, T. C. Steimle and M.C. Heaven, J. Chem. Phys.,134 031102 2011.